# Assembly Theory is an approximation to algorithmic complexity based on LZ compression that does not explain selection or evolution

### v5

@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
A critical examination of the AT methodology reveals significant shortcomings. Firstly, proponents of AT failed to conduct basic control experiments, a foundational aspect of introducing a new scientific metric. Benchmarking against established indices, particularly in coding and compression algorithms, is crucial to validating any new metric in the domain. Previous work on AT has never included meaningful experimental comparisons of the assembly index with other existing measures but schemes and mistaken theoretical arguments [8] (Figs. 1 and 2). Yet, we have shown that other algorithms, such as RLE, Huffman coding (one of the first, or the first dictionary-based universal compression algorithm), and other compression algorithms based on dictionary-based methods Shannon Entropy, reproduced equivalent or superior results compared to the results published by the AT [4].
@@
AT also introduces a cutoff value in its index, purported to offer a unique perspective on molecular complexity, distinct from those based on algorithmic complexity. This value indicates when an object is more likely to be organic, alive, or a product of a living system. However, various indices and compression algorithms tested have yielded equivalent cutoff values (see [4]). In fact, such an Assembly Index threshold replicates previous results on molecular separation that employ algorithmic complexity. Thus, this overlap in results challenges the notion that AT provides a unique tool for distinguishing between organic and non-organic matter. For instance, similar cutoff values have been derived in other studies, such as those mentioned in [5], effectively separating organic from non-organic molecular compounds. While the authors of AT ignored decades of research in chemical complexity based on graph theory and the principles of Entropy [13, 12, 37, 14], all of which are equivalent to AT, the depth of the purported experimental validation of AT was notably limited, especially when contrasted with more comprehensive studies. For instance, a more exhaustive and systematic approach, with proper control experiments involving over 15,000 chemical compounds [5] and employing algorithms from the LZ family (or Entropy) and others of greater sophistication (BDM), demonstrated the ability to separate organic from non-organic molecular compounds.
@@
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations in Fig. 1B, for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and the algorithm for copy number instantiated by the assembly index, is a restricted version of a Turing machine or finite automaton. Moreover, a Turing machine is a synonym of an algorithm and vice versa. It is a mistake to think of a Turing machine as a physical object or an object with particular properties or features (such as a head or a tape). All the algorithms in assembly theory are Turing machines and viceversa.
@@
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in AT [31]. This Venn diagram shows how AT is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf. A). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [38]) with an empty initial condition found by using a computable method (e.g. CTM, [39]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.
@@ 2.5 A circular argument cannot unify physics and biology
Central to Assembly Theory (AT) and the public claims made through their University Press Releases [9] is the connection that the authors of AT want to make to selection and evolution, maintaining that AT unifies physics and biology and explains and quantifies selection and evolution, both Darwinian and abiotic [2]. To demonstrate this connection, by assuming selectivity in the combination of linear strings $(P)$, the authors compare two schemes for combining linear strings $Q$, one random versus a non-random selection. This experiment yields observations of differences between the two $(R)$, and the authors conclude that selectivity $(S)$ exists in molecular assembly and, therefore, that AT can explain it.
@@
Yet, assuming selectivity in the combination of strings and proceeding to say that a selection algorithm is expected to differ from a random one makes for a circular argument. Furthermore, the conclusion $(S)$ reaffirms the initial assumption $(P)$, resulting in a circular argument. In a formal propositional chain, it can be represented as $P\implies Q\implies R\implies S\implies P$, where the conclusion merely restates the initial assumption, lacking any validation or verification, empirical or logical beyond a self-evident tautology.
@@
Nevertheless, the question of random versus non-random selection and evolution in the context of approximations to algorithmic complexity, including copy counting, was experimentally tested, empirically supported, and reported before in [6] following basic literature search and review, control experiments (comparison to other measures), and validation against existing knowledge in genetics and cell biology.
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. The suggestion is a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time, however, has always been fundamental in evolutionary theory, and claims about the object history are the foundations of evolutionary theory. AT revisits existing complexity science concepts described in the timeline provided in Fig. 2 but without acknowledgment or attribution.
Figure 2: A timeline of results in complexity science relevant to the claims and results from AT. AT renames several concepts such as dictionary trees as ‘assembly pathways’, heavily uses algorithmic probability in their reduction of combinatorial space arguments without attribution, and, as demonstrated, the assembly index is a LZ compression scheme (proofs provided in the Sup. Inf. A)
@@ 3 Conclusions
In this paper, we have shown that the calculation of the assembly index value from the minimum rooted assembly (sub)space is a compression scheme belonging to the LZ family of compression algorithms, rather than only being similar. Diving into the details of their algorithms and methods, we have revealed that they are equivalent to LZW’s and, therefore, to (Shannon) Entropy-based methods.
Our results demonstrate that an object with a low assembly index has high LZ compressibility (i.e., it is more compressible according to the LZ scheme) and, therefore, would necessarily display low entropy when generated by i.i.d. stochastic processes (or low entropy rate in ergodic stationary processes in general). In the opposite direction, an object with a high assembly index will have low LZ compressibility and, therefore, high entropy. This means that the methods based on AT, Shannon Entropy, and LZ compression algorithms are indistinguishable from each other with regard to the quantification of complexity of the assembling process, and the claim that those other approaches are incapable of dealing with their type of data (capturing ‘structure’ or anything that could not have been already characterised by Shannon Entropy or LZW, as it has been done before) is inaccurate.
@@
In addition, both theoretical concepts and empirical results in [1, 41, 8, 2] have been reported previously in earlier work in the context of AIT (from which algorithmic complexity is one of its indexes) for chemical processes in [5], and biology in [5, 16, 6, 17] but with comparison to other measures, including algorithms that go beyond statistical compression and connected to the concept of causal discovery.
Algorithmic complexity has been shown to be profoundly connected to causality [42, 43, 29, 11], not only for strings, but also it has been applied to images [17], networks [10, 44, 45, 46], vectors, matrices, and tensors in multiple dimensions [47, 48, 49]; and chemical structures [5]. In contrast to statistical compression schemes on which AT is based, whose assembly index method is a particular case of compression scheme, exploring other resource-bounded computable approximations to algorithmic complexity [26, 29], that consider aspects other than statistics, such as identical repetitions, may have more discriminatory power. This is because they may tell apart cases in which a high assembly index value with an expected low copy number (or frequency of occurrence in the ensemble), low LZ compressibility and high entropy, would indicate that an object is more causally disconnected, independent, or statistically random, but actually is causally dependent due to being more compressible by exploiting, for example, short rewriting rules describing the object.
One positive aspect of AT is that it offers a graph-like representation approach (albeit more convoluted than necessary) to approximate LZ compression in the specific context of molecular complexity, potentially making it more accessible to a broader and less technical community. However, it is inaccurate and a bad practice to ignore, omit or imply the absence of previous work (e.g. [5, 6, 7]), belittle and devalue the theories and ideas upon which their indexes and measures are entirely dependable and connected with (Shannon Entropy, compression, and Kolmogorov complexity), and make disproportionate public claims related to the contributions of AT (even calling it a theory that ‘unifies biology and physics’ and a ‘theory of everything’ by the media and the authors).
@@
We argue that the results and claims made by AT, in fact, highlight a working hypothesis that information and computation underpin the concepts necessary to explain the processes and building blocks of physical and living systems, concepts which the methods and frameworks of AT have been heavily inspired upon [51, 52].
@@ A.1 Mathematical framework
As in [53], let $\left(\Gamma,\phi\right)$ denote either an assembly space or an assembly subspace. In the interests of of simplifying notation, one can refer to $\left(\Gamma,\phi\right)$ simply as $\Gamma$ where the edge-labellings given by $\phi$ are not relevant. From [53], we have it that $c_{\Gamma}\left(x\right)$ denotes the assembly index of the object $x$ in the assembly space $\Gamma$.
Let $\mathcal{S}=\left(\mathbf{\Gamma},\mathbf{\Phi},\mathcal{F}\right)$ be an infinite assembly space, where every assembly space $\Gamma\in\mathbf{\Gamma}$ is finite, $\mathbf{\Phi}$ is the set of the correspondent edge-labelling maps $\phi_{\Gamma}$ of each $\Gamma$, and $\mathcal{F}=\left(f_{1},\dots,f_{n},\dots\right)$ is the infinite sequence of embeddings (in which each embedding is also an assembly map as in [53]) that ends up generating $\mathcal{S}$. That is, each $\,f_{i}\colon\left\{\Gamma_{i}\right\}\subseteq\mathbf{\Gamma}\to\left\{\Gamma_{i+1}\right\}\subseteq\mathbf{\Gamma}\,$ is a particular type of assembly map that embeds a single assembly subspace into a larger assembly subspace so that the resulting sequence of nested assembly subspaces defines a total order $\preceq_{\mathcal{S}}$, where
@@
In other words, $\mathcal{S}$ is the nested family of all possible finite assembly spaces from the same basis set (i.e., the root vertex that represents the set of all basic building blocks) such that $\mathcal{S}$ is infinite and enumerable by a program. [^2] Notice that an alternative proof of the results presented in this paper can be achieved just with the (more general) assumption that the assembly space may be finite but only needs to be sufficiently large in comparison to the deceiving program in Lemma 1.
@@
According to the assumptions and claims in [1, 53, 2] (as also discussed in [4]), notice that the values of $N_{T}$, $N$, $n_{i}$, $a_{i}$ are retrievable by an effective procedure from a given (finite) ensemble $X$, computational methods which are proposed by Assembly Theory [1, 53, 2]. Thus, function $A\#$ is a recursive function (i.e., a function that can be calculated by an algorithm) given a finite ensemble $X$, which in turn the authors in [1, 53, 2] claim to be one of the advantages of Assembly Theory with respect to algorithmic complexity.
@@
We have it that the (prefix) algorithmic complexity, denoted by $\mathbf{K}\left(x\right)$, is the length of the shortest prefix-free (or self-delimiting) program that outputs the encoded object $x$ when this program is run according to a (universal) programming language.
In accordance with the claims in [1, 41, 53, 2, 8], notice that the computability and feasibility of Assembly Theory’s methods directly imply the existence of algorithms that can enumerate the assembly spaces. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that Assembly Theory arbitrarily chooses to be based on, one has it that for any particular pathway assembly resulting in a molecule, there is a correspondent algorithm for which one can apply the methods described in [1, 41, 53] to calculate the pathway probability, and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is recursive from a given assembly space, which is already an assumed condition in [1, 41, 53]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run. For every allowed generative process that can assemble objects and thereby build another object, there is an algorithm that performs this process. Conversely, for every object, one also has it that there is a corresponding generative process allowed by Assembly Theory, a process which is effected by an algorithm. Therefore, according to the assumptions in [1, 41, 53, 2, 8], there is a formal theory $\mathbf{F}$ that contains Assembly Theory and all the procedures effected by algorithms, including the chosen method for calculating the assembly index of an object, the program that decides whether or not the criteria for building the assembly spaces are met, and so on.

### v6

@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
A critical examination of the AT methodology reveals significant shortcomings. Firstly, proponents of AT failed to conduct basic control experiments, a foundational aspect of introducing a new scientific metric. Benchmarking against established indices, particularly in coding and compression algorithms, is crucial to validating any new metric in the domain. Previous work on AT has never included meaningful experimental comparisons of the assembly index with other existing measures on false grounds that their measure is completely different [8] (Figs. 1 and 2). Yet, we have shown that other algorithms, such as RLE, Huffman coding (the first dictionary-based universal compression algorithm), and other compression algorithms based on dictionary-based methods and Shannon Entropy produced equivalent or superior results compared to the results published by the authors of AT [4].
@@
AT also introduces a cutoff value in its index, purported to offer a unique perspective on molecular complexity, distinct from those based on algorithmic complexity. This value indicates when an object is more likely to be organic, alive, or a product of a living system. However, various indices and compression algorithms tested have yielded equivalent cutoff values (see [4]). In fact, such an Assembly Index threshold replicates previous results on molecular separation that employ algorithmic complexity. Thus, this overlap in results challenges the notion that AT provides a unique tool for distinguishing between organic and non-organic matter. For instance, similar cutoff values have been derived in other studies, such as those mentioned in [5], effectively separating organic from non-organic molecular compounds. While the authors of AT ignored decades of research in chemical complexity based on graph theory and the principles of Entropy [13, 12, 38, 14], all of which are equivalent to AT, the depth of the purported experimental validation of AT has been notably limited, especially when contrasted with more comprehensive studies. For instance, a more exhaustive and systematic approach, with proper control experiments involving over 15,000 chemical compounds [5] and employing algorithms from the LZ family (or Entropy) and others of greater sophistication (BDM), demonstrated the ability to separate organic from non-organic molecular compounds.
@@
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations in Fig. 1B, for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and the algorithm for copy number instantiated by the assembly index represent a restricted version of a Turing machine or finite automaton. Moreover, a Turing machine is a synonym of an algorithm and vice versa. It is a mistake to think of a Turing machine as a physical object or an object with particular properties or features (such as a head or a tape). All the algorithms in assembly theory are Turing machines, and vice versa.
@@
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in AT [31]. This Venn diagram shows how AT is connected to and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf. A). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [39]) with an empty initial condition found by using a computable method (e.g. CTM, [40]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.
@@ 2.5 A circular argument cannot unify physics and biology
Central to Assembly Theory (AT) and the public claims made through the authors’ university press releases [9] is the connection that they make to selection and evolution, maintaining that AT unifies physics and biology and explains and quantifies selection and evolution, both Darwinian and abiotic [2]. To demonstrate this connection, assuming selectivity in the combination of linear strings $(P)$, the authors compare two schemes for combining linear strings $Q$, one random versus a non-random selection. This experiment yields observations of differences between the two $(R)$, and the authors conclude that selectivity $(S)$ exists in molecular assembly, and therefore that AT can explain it.
@@
Yet, assuming selectivity in the combination of strings and proceeding to say that a selection algorithm is expected to differ from a random one makes for a circular argument. Furthermore, the conclusion $(S)$ reaffirms the initial assumption $(P)$, resulting in a circular argument. In a formal propositional chain, it can be represented as $P\implies Q\implies R\implies S\implies P$, where the conclusion merely restates the initial assumption, lacking any validation or verification, empirical or logical, beyond a self-evident tautology.
@@
Nevertheless, the question of random versus non-random selection and evolution in the context of approximations to algorithmic complexity, including copy counting, was experimentally tested, empirically supported, and reported before in [6] following proper basic principles such as a literature search and review, control experiments (comparison to other measures), and validation against existing knowledge in genetics and cell biology.
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. The suggestion is a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time, however, has always been fundamental in evolutionary theory, and claims about an object’s history are the foundations of evolutionary theory. AT revisits existing complexity science concepts described in the timeline provided in Fig. 2 but without acknowledgment or attribution.
Figure 2: A timeline of results in complexity science relevant to the claims and results of AT. AT renames several concepts, e.g. dictionary trees as ‘assembly pathways’; relies heavily on algorithmic probability in its reduction of combinatorial space arguments, without attribution; and, as demonstrated, the assembly index is a LZ compression scheme (proofs provided in the Sup. Inf. A)
@@ 3 Conclusions
In this paper, we have shown that the calculation of the assembly index value from the minimum rooted assembly (sub)space is a compression scheme belonging to the LZ family of compression algorithms, rather than merely being similar. Diving into the details of AT’s algorithms and methods, we have revealed that they are equivalent to LZW’s, and therefore to (Shannon) Entropy-based methods.
Our results demonstrate that an object with a low assembly index has high LZ compressibility (i.e., it is more compressible according to the LZ scheme), and therefore would necessarily display low entropy when generated by i.i.d. stochastic processes (or low entropy rate in ergodic stationary processes in general). In the opposite direction, an object with a high assembly index will have low LZ compressibility, and therefore high entropy. This means that the methods based on AT, Shannon Entropy, and LZ compression algorithms are indistinguishable from each other with regard to the quantification of complexity of the assembling process, and the claim that these other approaches are incapable of dealing with AT’s type of data (capturing ‘structure’ or anything that could supposedly not be characterised by Shannon Entropy or LZW) is inaccurate.
@@
In addition, both theoretical concepts and empirical results in [1, 42, 8, 2] have been reported previously in earlier work in relation to chemical processes in [5], and to biology in [5, 16, 6, 17] but with comparison to other measures, including methods that go beyond traditional statistical compression and are connected to the concept of causal discovery [10, 29].
Algorithmic complexity has been shown to be profoundly connected to causality [43, 44, 29, 11], and not only in strings. It has also been applied to images [17], networks [10, 45, 46, 47], vectors, matrices, and tensors in multiple dimensions [48, 49, 50]; and to chemical structures [5]. In contrast to the statistical compression schemes on which AT is based, and whose assembly index method is a particular case of a compression scheme, exploring other resource-bounded computable approximations to algorithmic complexity [26, 29] that consider aspects other than traditional statistical patterns such as identical repetitions, may have more discriminatory power. This is because they may tell apart cases with a high assembly index value and an expectedly low copy number (or frequency of occurrence in the ensemble) from those with low LZ compressibility and high entropy, seeming to indicate that an object is more causally disconnected, independent, or statistically random, while actually, it is strongly causally dependent due to non-trivial rewriting rules not captured by statistical means. This is because AT is to statistical correlation, as measures of algorithmic complexity are to causation. Rarely the two are the same (correlation is not causation) and when they are, they are fully captured by other traditional statistical measures such as Shannon Entropy without the introduction of a methodological different framework that adds no more to the correlation problem beyond Shannon Entropy.
If AT has some elements that appear to have discriminatory power [42, 1], it is because of its connection to Shannon Entropy.
One positive aspect of AT is that it offers a graph-like representational approach (albeit more convoluted than necessary) to approximating LZ compression in the specific context of molecular complexity, potentially making it more accessible to a broader and less technically-minded community. However, it is inaccurate and bad practice to ignore, omit or imply the absence of previous work (e.g. [5, 6, 7, 32]), belittle and devalue the theories and ideas which AT’s indexes and measures are entirely dependent upon and connected to (Shannon Entropy, compression, and Kolmogorov complexity), and make disproportionate public claims related to the contributions of AT (with the media and the authors themselves going so far as to call it a theory that ‘unifies biology and physics’ and a ‘theory of everything’).
@@
We argue that the results and claims made by AT, in fact, highlight a working hypothesis that information and computation underpin the concepts necessary to explain the processes and building blocks of physical and living systems, concepts which the methods and frameworks of AT have been heavily inspired by [52, 53].
@@ A.1 Mathematical framework
As in [54], let $\left(\Gamma,\phi\right)$ denote either an assembly space or an assembly subspace. In the interests of simplifying notation, one can refer to $\left(\Gamma,\phi\right)$ simply as $\Gamma$ where the edge-labellings given by $\phi$ are not relevant. From [54], we have it that $c_{\Gamma}\left(x\right)$ denotes the assembly index of the object $x$ in the assembly space $\Gamma$.
Let $\mathcal{S}=\left(\mathbf{\Gamma},\mathbf{\Phi},\mathcal{F}\right)$ be an infinite assembly space, where every assembly space $\Gamma\in\mathbf{\Gamma}$ is finite, $\mathbf{\Phi}$ is the set of the corresponding edge-labelling maps $\phi_{\Gamma}$ of each $\Gamma$, and $\mathcal{F}=\left(f_{1},\dots,f_{n},\dots\right)$ is the infinite sequence of embeddings (in which each embedding is also an assembly map as in [54]) that ends up generating $\mathcal{S}$. That is, each $\,f_{i}\colon\left\{\Gamma_{i}\right\}\subseteq\mathbf{\Gamma}\to\left\{\Gamma_{i+1}\right\}\subseteq\mathbf{\Gamma}\,$ is a particular type of assembly map that embeds a single assembly subspace into a larger assembly subspace so that the resulting sequence of nested assembly subspaces defines a total order $\preceq_{\mathcal{S}}$, where
@@
In other words, $\mathcal{S}$ is the nested family of all possible finite assembly spaces from the same basis set (i.e., the root vertex that represents the set of all basic building blocks) such that $\mathcal{S}$ is infinite and enumerable by a program. [^2] Note that an alternative proof of the results presented in this paper can be achieved just with the (more general) assumption that the assembly space may be finite but only needs to be sufficiently large in comparison to the deceiving program in Lemma 1.
@@
According to the assumptions and claims in [1, 54, 2] (as also discussed in [4]), notice that the values of $N_{T}$, $N$, $n_{i}$, $a_{i}$ are retrievable by an effective procedure from a given (finite) ensemble $X$, computational methods which are proposed by Assembly Theory [1, 54, 2]. Thus, function $A\#$ is a recursive function (i.e., a function that an algorithm can calculate) given a finite ensemble $X$, a property which the authors of [1, 54, 2]falsely claim to be one of the advantages of Assembly Theory over algorithmic complexity, because this template was computable already to begin with, so there was no room for uncomputability claims.
@@
We have it that the (prefix) algorithmic complexity, denoted by $\mathbf{K}\left(x\right)$, is the length of the shortest prefix-free (or self-delimiting) program that outputs the encoded object $x$ when this program is run using a (universal) programming language.
In accordance with the claims made in [1, 42, 54, 2, 8], note that the computability and feasibility of Assembly Theory’s methods directly imply the existence of algorithms that can enumerate the assembly spaces. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that Assembly Theory arbitrarily chooses to be based on, one has it that for any particular pathway assembly resulting in a molecule, there is a corresponding algorithm for which one can apply the methods described in [1, 42, 54] to calculate the pathway probability, and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is recursive from a given assembly space, which is already an assumed condition in [1, 42, 54]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run. For every permitted generative process that can assemble objects and thereby build another object, there is an algorithm that performs this process. Conversely, for every object, one has it that there is a corresponding generative process allowed by Assembly Theory, a process which is effected by an algorithm. Therefore, according to the assumptions in [1, 42, 54, 2, 8], there is a formal theory $\mathbf{F}$ that contains Assembly Theory and all the procedures effected by algorithms, including the chosen method for calculating the assembly index of an object, the program that decides whether or not the criteria for building the assembly spaces are met, and so on.